alcon Crest Aces (FCA), Inc., is considering the purchase of a small plane to use in its wing-walking demonstrations and aerial tour business. Various information about the proposed investment follows:     

Initial investment $ 280,000
Useful life $ 10 years
Salvage value 25,000
Annual net income generated $ 6,200
FCA's cost of capital 8 %

1. Accounting rate of return. (Round your answer to 2 decimal places.)

2. Payback period. (Round your answer to 2 decimal places.) (Find how many years)

3. Net present value (NPV)

4. Recalculate FCA's NPV assuming the cost of capital is 3% percent.

5. Without doing any calculations, what is the project's IRR?
Less than 3%
Between 3% and 8%

Greater than 8%
Answer #1

1.

Average annual income = $6,200

Average investment = 1/2 x (Initial investment - Scrap value)

= 1/2 x (280,000 - 25,000)

= 1/2 x 255,000

= $127,500

Accounting rate of return = Average annual profit/Average investment

= 6,200/127,500

= 4.86%

2.

Annual depreciation = (Initial investment - Scrap value)/Useful life

= (280,000 - 25,000)/10

= $25,500

Annual cash inflow = Annual income + Depreciation

= 6,200 + 25,500

= $31,700

Payback period = Initial investment/Annual cash inflow

= 280,000/31,700

= 8.83 years

3.

NPV = Present value of cash inflows - Present value of cash outflows

= Annual cash inflow x PVAF(8%, 10) + Terminal cash inflow x PVF(8%, 10) - 280,000

= 31,700 x 6.710 + 25,000 x 0.463 - 280,000

= 212,707 + 11,575 - 280,000

= - $55,718

4.

NPV = Present value of cash inflows - Present value of cash outflows

= Annual cash inflow x PVAF(3%, 10) + Terminal cash inflow x PVF(3%, 10) - 280,000

= 31,700 x 8.530 + 25,000 x 0.744 - 280,000

= 270,401 + 18,600 - 280,000

= $9,001

5.

Since at 8% discount rate, NPV is negative and at 3% discount rate, NPV is positive, hence IRR must lie between 3% to 8%.